Given α is non-income tax, β is income tax, δ is marginal propensity to consume, γ is autonomous consumption, T (tax), Y (national income), I0 (investment) and G0 (government expenditure).


a) Formulate the equations needed to find the reduced form of equilibrium income (Ye).
1
Expert's answer
2020-11-17T07:25:04-0500

"Ye=C+I+G"

"C=Ca+MRC\times Yd"

"Yd=Ye-T+TR"

"T=Ta+t\times Ye"

substitute these designations into the equation

"T=α+β\times Ye"

"Yd=Ye-T+TR=Ye-(α+β\times Ye)"

"C=Ca+MRC\times Yd=γ+δ((Ye-(α+β\times Ye))"

"Ye=C+I+G=γ+δ((Ye-(α+β\times Ye))+I0 +G0"

"Ye=γ+δ((Ye-(α+β\times Ye))+I0 +G0"